# Network Intrusion Datasets: A Survey, Limitations, and Recommendations

### v1

@@ 6.2.1 Preparation
The preparation phase encompasses all activities prior to the actual data collection. Its steps include a) defining the dataset’s objective, b) designing relevant scenarios, and c) preparing the capture environment.
@@
Objective definition: The first step is to define the purpose of the dataset, which then guides its entire creation process – selecting scenarios, features, and post-processing. Since many different attacks and network topologies exist, creating a universal “perfect” dataset is infeasible. Instead, data authors should narrow down to the specific problem the dataset targets. Its objective should cover at least three aspects: 1) malicious traffic type, 2) network size, and 3) time span. For instance, “to benchmark NIDSs for Slow DoS detection on medium-sized enterprise networks over several weeks. A clear objective helps to determine the dataset’s limitations and boundaries.
@@
Scenarios design: Following the specified objective, the next step is to design scenarios simulated during the execution and collection step. Primary considerations include planning on obtaining normal background traffic, malicious traffic, and interactions between the two. The design phase also schedules network events – e.g., specific attacks, exact commands to generate them, and a relative timestamp of their start and ending. The dataset should reflect current traffic patterns and attacks in real-world operational networks, so recency and realism should be considered the top priorities. Designs are also generally advised to support updates, allowing new scenarios and traffic patterns to be incorporated on demand [31, 22].
@@
As discussed in Section 6.1, documentation is crucial for assessing data realism. Nevertheless, realism is primarily shaped during scenario design and environment setup, further influenced by nearly every other action during the dataset creation. Regarding design, data authors should create scenarios aligned with the patterns observable in real-world networks. To prevent spatial bias, datasets can also reflect real traffic ratios, typically with a significant disbalance favoring background traffic.
@@
Further considerations should be given to traffic diversity, ensuring the dataset includes diverse and variable normal and malicious traffic patterns. Insufficient variability can make the distinction between normal and malicious traffic trivial, leading to unrealistically overconfident NIDS performance. Extending capture duration by covering day-and-night or workweek-weekend cycles, such as in UGR'16 [104], is one way to enhance traffic variability. Although emulation and synthetic traffic generation methods are becoming more sophisticated, some argue that real background traffic needs to be used to achieve realism [13, 104]. However, it presents challenges like anonymization and establishing reliable ground truth [43].
@@
Environment setup: The last preparation step includes setting up a capture environment that aligns with the defined objectives and scenarios. In emulated setups, this includes configuring network equipment, end-hosts, as well as capture infrastructure, such as port mirroring or NetFlow probes, with realistic configuration and ensuring the capture will not bias the data (e.g., port mirroring introducing significant delays). Real environments utilize existing infrastructure but might add specific monitoring devices, vulnerable servers, or malicious hosts just for the capture. For synthetic data, the setup refers to training or configuring a model to generate the traffic.
@@ 6.2.2 Data Collection
Collecting the data comprises a) executing scenarios and collecting the traffic, and b) extracting relevant features.
Execution & collection: After setting up the environment, the scenarios are run according to the execution plan, and relevant traffic is captured. The data format depends on the dataset goals, but multiple data formats (and sources), such as packets, flows, and logs, are generally preferred to support benchmarking of various NIDS types [144]. However, collecting data from multiple sources presents challenges like clock synchronization and the need for event correlation.
Feature extraction: Feature extraction now processes the collected data into feature vectors suitable for intrusion detection. When creating a raw packet dataset or extracting features directly from NetFlow, this step is implicit during execution & collection. Otherwise, a relevant feature set needs to be selected in accordance with the objective of the dataset, i.e., effectively describing attacks while distinguishing them from normal background traffic. Despite efforts to standardize NID feature sets [40], the optimal feature set remains an open issue [201]. We thus encourage researchers to experiment with feature sets while ensuring they are a) practicalfor real-time detection (computable online) and b) accompanied by the extraction source code and raw data to facilitate reproducibility.
@@ 6.2.3 Post-Processing
The post-processing phase modifies the collected data to accommodate its sound evaluation. It involves three steps, namely data: a) labeling, b) validation, and c) cleaning.
Labeling: Full and direct labeling is a crucial property of NID benchmark datasets, ensuring that traffic classes are distinguishable via a simple look-up or filter upon the data. Although log files benefit certain NIDS types, they should not be used for indirect labeling due to introduced uncertainties and added effort in determining the labels. Instead, logs can supplement existing direct labels. Depending on the dataset setup, labeling can be automated or human-guided. Guerra et al. [206] provide an overview of NID labeling techniques.
Data validation: Validating the data ensures correctness and realism – key factors in quality assurance. Data correctness refers to error-free capture and alignment with the designed scenarios. This involves investigating for packet losses or malformations, verifying proper attack execution and its desired effects (e.g., DoS attack causing a service disruption), correct feature extraction (e.g., no unexpected N/A values, empty columns), the post-processing not introducing any artifacts or biases, and correct labeling. These checks are typically performed manually based on the data’s nature and semantics. In addition to verifying the design and capture process, realism might be validated using statistical tests, e.g., as presented by Schoen et al. [52], along with domain-knowledge checks.
Data cleaning: In NID data creation, cleaning mainly involves anonymization. Datasets collected from real networks must consider ethical, privacy, and legal issues – anonymizing IP addresses and packet payloads thus becomes necessary. Sometimes, more information like DNS queries, timestamps, or transport layer ports might need to be stripped. Nevertheless, excessive anonymization reduces realism, limiting data value [204]. Dropping entire entries might even be necessary if flagged as undesirable or erroneous. Nevertheless, we advocate for minimal changes to avoid the degradation of realism.
@@ 6.2.4 Publication
  • Goal of the dataset: Types of NIDS is the dataset for. Attacks and scenarios captured within it.
  • Data generation: Description of procedures to generate background traffic and attacks. List concrete tools, algorithms, and their parameters.
@@
The above 10-step process produces a static dataset reflecting network behavior within a specific period of time. However, as discussed in this paper, static datasets can quickly become outdated. For this reason, it is desirable to continually update datasets with new scenarios and traffic patterns by re-iterating these 10 steps or designing specialized methodologies for collecting continual datasets, as elaborated in Section 9.
@@ 6.3 Using a Dataset
After obtaining relevant data, the paper now briefly covers their proper handling. Similar to other domains, the key is to understand their semantics via exploratory data analysis and documentation, followed by standard ML procedures for data preparation like encoding and scaling. However, the following paragraphs present additional considerations specific to NID:
@@
You cannot see the future: Given the dynamic nature of network traffic (Section 3.1), benign and malicious traffic patterns can change over time. For this reason, it is crucial to respect the temporal properties of the data during evaluation – training data timestamps should strictly precede evaluation data to avoid temporal data snooping. Evaluation methods ignoring this aspect (including cross-validation) might introduce bias, leading to over-optimistic results [37].
@@
Some shortcuts lead to hell: Network data contain several features that might contaminate the process of data-driven learning (e.g., ML-based algorithms). These include IP addresses, port numbers, timestamps, flow IDs, or even Time to live (TTL) values. If used for training, the learner might spuriously correlate these artifacts with a specific activity, causing a shortcut learning phenomenon [208] and overestimating the model’s performance. For instance, D’Hooge et al. [209] demonstrated 70% – 100% accuracy on popular NID datasets by considering the destination port as the sole feature.
@@
For this reason, possible artifacts should be removed or encoded into general features (e.g., classifying IP addresses as local, global, or other) to preserve some of their informational value while preventing spurious correlations. However, encoding must be done cautiously, and the presence of spurious correlations should be verified. Identification of potential artifacts is task-specific, requiring domain knowledge and data understanding. Nevertheless, the features mentioned above ought to be problematic in most datasets and should be addressed.
@@
The heaviest hammer is not always the best: As shown by numerous studies [22, 210, 211], even simplistic tree-based models like Random Forest or Gradient Boosting achieve near state-of-the-art performance with significantly faster execution. D’Hooge et al. demonstrated that even single-rule supervised ML models perform well on most cybersecurity datasets [212] due to the over-correlation of features with labels [11]. For these reasons, we recommend starting with simple models before employing deep learning. For practical real-time intrusion detection, it is desirable to keep models simple so their inference (i.e., attack detection) is quick, enabling efficient processing of large traffic volumes. Therefore, we suggest adopting a data-centric perspective [213], focusing on understanding and enhancing the data with simple, explainable models rather than maximizing a specific metric with a complex black-box model without understanding the data.
@@
Once? Not enough!: Drawing conclusions based on a single evaluation run on a single dataset is generally insufficient yet common in the literature. Ideally, models should be evaluated on multiple datasets with various splits. Multiple runs (respecting temporal dependencies) provide results comparable with other solutions in a statistically significant manner via methods like the t-test [214] or the Mann-Whitney U-test [215].
Mix them with caution: Same as with alcohol, you should also think twice before mixing different datasets. Since their data come from different networks and time periods, their traffic patterns will likely differ. Therefore, combining datasets is incorrect by design since bias would probably be introduced. This can cause automated learners to focus on dataset-specific properties (e.g., inter-packet arrival times) rather than malicious patterns as intended. Therefore, we strongly advise against mixing data from different sources unless necessary.
@@
Tools for network traffic replay like tcpreplay [217] preserve packet contents while unifying time dependencies within a single environment, thus reducing time-related biases. However, merging distinct datasets can still introduce artifacts. For instance, data from different systems (e.g., Linux vs. Windows) might present differences in packet contents, such as the employed encryption protocols. Additionally, replaying complex attacks requiring both sides of the communication might be challenging, as most replay tools are stateless [183]. These factors can introduce additional hidden biases during data merging. For this reason, NIDS methods should also be benchmarked on other datasets originating from a single source to prevent potential biases from significantly skewing the results.
These paragraphs presented key considerations for NIDS evaluations. For a more in-depth discussion, we recommend the articles by Arp et al. [14] and Apruzzese et al. [4].

### v2

@@ 6.2.1 Preparation
The preparation phase encompasses all activities prior to the actual data collection. Its steps involve a) defining the dataset’s objective, b) designing relevant scenarios, and c) preparing the capture environment.
  1. 1.
@@
    Objective definition: The first step is to define the purpose of the dataset, which then guides its entire creation process – selecting scenarios, features, and post-processing. Since many different attacks and network environments exist, creating a universal “perfect” dataset is infeasible. Instead, dataset authors should narrow down to the specific problem (e.g., attack class or adversary tactics). Its objective should cover at least three aspects: 1) malicious traffic class(es), 2) environment type and its size, and 3) time span. For instance, “to benchmark NIDSs for Slow DoS detection on medium-sized enterprise networks over several weeks. A clear objective helps to determine the dataset’s scope and delimitations.
  2. 2.
@@
    Scenario design: Following the specified objective, the next step is to design the scenarios simulated during the execution and collection. Key considerations include planning on obtaining normal background traffic, malicious traffic, and their interactions. This phase also schedules network events – e.g., specific attacks, exact commands to generate them, and their relative start and end timestamps. The dataset should reflect current traffic patterns and attacks in real-world operational networks, making timeliness and realism top priorities. Designs are also advised to be extensible, allowing new scenarios and traffic patterns to be incorporated on demand [32, 23].
@@
    As discussed in Section 6.1, documentation is crucial for assessing data realism. Nevertheless, realism is primarily shaped during scenario design and environment setup, further influenced by nearly every other action during the dataset creation. Regarding design, data authors should create scenarios aligned with the patterns observable in real-world networks. To prevent spatial bias [39], datasets can also reflect realistic traffic ratios, typically by a disbalance favoring background traffic.
@@
    Further considerations should be given to diversity, ensuring the dataset includes diverse and variable normal and malicious traffic profiles. Insufficient variability can make the distinction between normal and malicious traffic trivial, leading to unrealistically overconfident NIDS performance. Extending capture duration by covering day-and-night or workweek-weekend cycles, such as in UGR'16 [105], is one way to enhance traffic variability. Although emulation and synthetic traffic generation methods are becoming more sophisticated, some argue that real background traffic needs to be used to achieve realism [14, 105]. However, it presents challenges like privacy preservation and establishing reliable ground truth [45].
  3. 3.
@@
    Environment setup: The last preparation step includes setting up a capture environment that aligns with the defined objectives and scenarios. In emulated setups, this includes configuring network equipment, end-hosts, as well as capture infrastructure, such as port mirroring or NetFlow probes, with realistic configuration and ensuring the capture will not bias the data (e.g., port mirroring introducing significant delays). Real environments utilize existing infrastructure but might add specific monitoring devices, vulnerable servers, or malicious hosts just for simulation purposes. For synthetic data, this setup refers to training or configuring a model to generate the traffic.
@@ 6.2.2 Data Collection
Collecting the data comprises a) executing scenarios and collecting the traffic and b) extracting relevant features.
  1. 4.
    Execution & collection: After setting up the environment, the scenarios are run according to the execution plan, and relevant traffic is captured. The data format depends on the dataset goals, but multiple data formats (and sources), such as packets, flows, and logs, are generally preferred to support benchmarking of various NIDS types [144]. However, collecting data from multiple sources presents challenges, such as clock synchronization and the need for event correlation.
  2. 5.
    Feature extraction: Feature extraction now processes the collected data into feature vectors suitable for intrusion detection. When creating a raw packet dataset or extracting features directly from NetFlow, this step is implicit during execution and collection. Otherwise, a relevant feature set needs to be selected in accordance with the dataset’s objective, i.e., effectively describing attacks while distinguishing them from normal background traffic. Despite efforts to standardize NID feature sets [42], the optimal feature set remains an open issue [203]. We thus encourage researchers to experiment with feature sets while ensuring they are: a) practicalfor real-time detection (computable online), and b) accompanied by the extraction source code and raw data to facilitate reproducibility.
@@ 6.2.3 Post-Processing
The post-processing phase modifies the collected data to accommodate its usage as benchmarks. It involves three steps, namely data: a) labeling, b) validation, and c) cleaning.
  1. 6.
    Labeling: Complete and direct labeling is a crucial property of NID benchmark datasets, ensuring that traffic classes are distinguishable via a simple look-up or filter upon the data. Although log files benefit certain NIDS types, they should not be used for indirect labeling due to introduced uncertainties and added effort in determining the labels. Instead, logs can supplement existing direct labels. Depending on the dataset setup, labeling can be automated or human-guided. Guerra et al. [47] provide an overview of NID labeling techniques.
  2. 7.
    Data validation: Validating the data ensures correctness and realism – key factors in quality assurance. Data correctness refers to error-free capture and alignment with the designed scenarios. This involves investigating for packet losses or malformations, verifying proper attack execution and its desired effects (e.g., a DoS attack causing a service disruption), correct feature extraction (e.g., no unexpected N/A values, empty columns), the post-processing not introducing any artifacts or biases, and correct labeling. These checks are typically performed manually based on the data’s nature and semantics, although automated tests, as proposed by Flood et al. [8], are also viable. In addition to verifying data correctness, their realism can be validated using statistical tests along with domain-knowledge checks, as presented by Schoen et al. [56].
  3. 8.
    Data cleaning: In NID data creation, cleaning primarily refers to privacy preservation – i.e., anonymizing IP addresses and packet payloads to address ethical, privacy, and legal concerns for data collected from real-world networks. Additional information, like DNS queries, timestamps, or transport layer ports, may also need to be stripped. In some cases, removing entire erroneous or undesirable entries might even be necessary. However, excessive anonymization can reduce realism, limiting the dataset’s value [207]. For this reason, we advocate for minimal changes to avoid the degradation of realism.
@@ 6.2.4 Publication
    • Goal of the dataset: Types of NIDS is the dataset designed for. Attacks and scenarios captured within it.
    • Data generation: A description of the procedures to generate background traffic and attacks, including specific tools, algorithms, and their parameters.
@@
The above 10-step process produces a static dataset reflecting network behavior within a specific period of time. However, as discussed in this paper, static datasets can quickly become outdated. For this reason, it is desirable to continually update datasets with new scenarios and traffic patterns by re-iterating these steps or designing specialized methodologies for collecting continual datasets, as elaborated in Section 7.
@@ 6.3 Using a Dataset
After obtaining relevant data, the paper now briefly covers their proper handling. Similar to other domains, the key is to understand their semantics via exploratory data analysis and documentation, followed by standard ML procedures for data preparation like encoding and scaling [210]. However, additional NID-specific considerations should be made. While this subsection summarizes them only briefly, works by Arp et al. [15] and Apruzzese et al. [4] provide more detailed analysis.
@@
You cannot see the future: Given the dynamic nature of network traffic (Section 3.1), benign and malicious traffic patterns can change over time. For this reason, it is crucial to respect the temporal properties of the data during evaluation – training data timestamps should strictly precede evaluation data to avoid temporal data snooping. Evaluation methods ignoring this aspect (including cross-validation) might introduce bias, leading to over-optimistic results [15, 39].
@@
Some shortcuts lead to hell: Network data contain several features that might contaminate the process of data-driven learning (e.g., ML-based algorithms). These include IP addresses, port numbers, timestamps, flow IDs, or even Time to live (TTL) values. If used for training, the learner might spuriously correlate these artifacts with a specific activity, causing a shortcut learning phenomenon [211] and overestimating the model’s performance [15, 33]. For instance, D’Hooge et al. [212] demonstrated 70% – 100% accuracy on popular NID datasets by considering the destination port as the sole feature.
@@
For this reason, possible artifacts should be removed or encoded into general features (e.g., classifying IP addresses as local, global, or other) to preserve some of their informational value while preventing spurious correlations. However, encoding must be done cautiously, and the presence of spurious correlations should be verified. Identification of potential artifacts is task-specific, requiring domain knowledge and data understanding. Nevertheless, the features mentioned above ought to be problematic in most datasets and should be addressed. Heuristic identification of highly dependent features (potential artifacts) can be performed by training a model to distinguish between the attack and the background traffic using a single feature [8]. Flood et al. [8] suggest justifying the model’s performance via ML explainability techniques [213, 214] and connecting important features to attacks’ properties.
@@
The heaviest hammer is not always the best: As shown by numerous studies [23, 215, 216], even simplistic tree-based models like Random Forest or Gradient Boosting achieve near state-of-the-art performance with significantly faster execution. D’Hooge et al. demonstrated that even single-rule supervised ML models perform well on most cybersecurity datasets [217] due to the over-correlation of features with labels [12]. Therefore, a general recommendation is to start with simple models before employing deep learning [33]. For practical real-time intrusion detection, it is desirable to keep models simple so their inference (i.e., attack detection) is quick, enabling efficient processing of large traffic volumes. Therefore, we suggest adopting a data-centric perspective [218], focusing on understanding and enhancing the data with simple, explainable models rather than maximizing a specific metric with complex black-box models without understanding the data.
@@
Regarding specific metrics, Arp et al. [15] suggest selecting them based on practical domain needs – primarily false alarm and detection rates for NID. To tackle the base rate fallacy, it is beneficial to incorporate metrics that account for class imbalance, namely precision, recall, precision-recall curve, as well as the Receiver Operating Characteristics (ROC) curve and the area under it (AUC-ROC). Therefore, it is crucial to utilize multiple metrics rather than relying on a single one.
Once? Not enough! Drawing conclusions based on a single evaluation run on a single dataset is generally insufficient yet relatively common in the literature [4]. Ideally, models should be evaluated on multiple datasets with various splits. Conducting multiple runs (respecting temporal dependencies) allows for statistically significant comparisons using methods such as the t-test [219] or the Mann-Whitney U-test [220].
Mix them with caution: Same as with alcohol, you should also think twice before mixing different datasets. Since their data come from different networks and time periods, their traffic patterns will likely differ. Therefore, combining datasets is incorrect by design since bias would likely be introduced. This can cause automated learners to focus on dataset-specific properties (e.g., inter-packet arrival times) rather than malicious patterns as intended. Therefore, we strongly advise against mixing data from different sources unless necessary.
@@
Tools for network traffic replay like tcpreplay [222] preserve packet contents while unifying time dependencies within a single environment, thus reducing time-related biases. However, merging distinct datasets can still introduce artifacts. For instance, data from different systems (e.g., Linux vs. Windows) might present differences in packet contents, such as the employed encryption protocols. Additionally, replaying complex attacks requiring both sides of the communication might be challenging, as most replay tools are stateless [185]. These factors can introduce additional hidden biases during data merging. For this reason, NIDS methods should also be benchmarked on other datasets originating from a single source to prevent potential biases from skewing the results.
7 Future Domain Directions
7.1 Data Generation
The research branch of data generation focuses on generating timely and realistic data – either by collecting real-world traces or improving traffic emulation and synthesis techniques.
Collecting timely real-world data: One of the key NID domain limitations is the lack of realistic real-world samples. As our survey revealed, only one in five datasets contains real traffic, with this ratio decreasing in recent years. This limitation impacts NIDS generalization, leading to overfitting to unrealistic traffic patterns and reduced performance in operational environments. For this reason, more data with contemporary real-world network patterns are required. While datasets like CTU-13 [59], Appraise H2020 [156], and LITNET [128] incorporate real traffic, they remain static and thus quickly become outdated as network traffic patterns evolve.
Improving simulated normal traffic: Although obtaining real-world traffic is generally considered ideal, it is often difficult or impossible due to privacy concerns, the absence of desired attack events, or operational constraints. For this reason, realistic traffic emulation remains crucial for constructing diverse and representative NID datasets. While attack traffic can be emulated via existing tools used by real-world attackers, creating realistic benign traffic remains the primary challenge.
To enhance traffic realism and robustness, future research should focus on bridging the gap between simulated (emulated and synthetic) benign traffic and its real-world counterpart. Although models like Generative Adversarial Networks (GANs) [187] and Bayesian Networks (BNs) [56] have already been explored, investigating and comparing novel generative methods like diffusion models [227] for benign network traffic synthesis offer promising research directions for diverse, realistic, and privacy-preserving traffic creation. Furthermore, the synthetic traffic must respect real-world constraints, making it important to formalize domain-specific conditions for synthetic traffic validity (e.g., network flows cannot have zero bytes).
Despite their benefits, reproducibility and extensibility have received little attention from the community. Therefore, most virtualized testbed datasets lack environment setup files, configuration details, or information on customizing the environment. Although AIT-LDSv2.0 [45] emphasizes reproducibility, its environment customizability is limited. Future research should thus focus on proposing standardized methods, frameworks, and templates for fully replicable and extensible NID data creation in virtualized and cloud environments to enable dataset validation, modification, customization, and extension.
In addition to previous scenarios, Hindy et al. [46] and Keersmaeker et al. [229] pointed out that many attacks and IoT environments are not covered within public NID datasets at all. For these reasons, future research should prioritize covering specific scenarios and environments, along with ensuring robust coverage of emerging threats to streamline NIDS research in these areas. However, without continually captured and extensible datasets is made, new datasets will remain susceptible to getting outdated due to the ever-changing nature of NID.
7.2 Data Processing
Exploring data granularities for intrusion detection: As shown in our survey, most datasets are flow-based, while nearly half lack packet data. Flow-based NID has become the de facto standard for NID research. Similarly, many studies have explored the selection of optimal flow-based features [230, 231, 232], but little research has been done on their effectiveness across different intrusion detection scenarios.
7.3 Data Validation
Formalizing frameworks for quality evaluation: While some metrics for NID data quality evaluation exist [205, 104], they primarily serve to compare datasets rather than assess their intrinsic quality (i.e., realism, correctness, and usability). Furthermore, their interpretation is often ambiguous due to the lack of usage instructions. Additionally, Soukup et al. [206] employed completeness and reliability measures with an evolutionary algorithm to assess dataset quality. However, their study presents only preliminary results and lacks detailed information on the proposed framework, limiting its practical applicability. These deficiencies make evaluating and selecting suitable NID datasets difficult, leading to frequent utilization of flawed datasets that undermine research reliability.
A recent paper by Flood et al. [8] proposed several manual and automated tests to audit common limitations of NID data, such as mislabeling, artifacts, and lack of diversity. Although this represents a significant step toward data quality evaluation and the selection of suitable datasets, their heuristics focus only on a subset of typical problems, while the proposed manual analysis requires access to both flow-based and raw data.
In order to improve NID data evaluation, future research should formalize methods and develop frameworks for evaluating NID data validity and quality, i.e., ensure datasets are both realistic and error-free. Current data validation heavily relies on manual analysis, as seen in studies on CIC-IDS2017 and CSE-CIC-IDS2018 issues [108, 48, 165], with Engelen et al. [226] recommending manual analysis by domain experts before public data release. Such validations are time-consuming and dependent on the availability of experts, posing additional challenges for data creators. Therefore, research on NID data quality should also integrate expert knowledge into (semi-)automated validation methods, potentially building on the Flood et al.’s work [8]. Standardized approaches to verify other aspects of data quality – e.g., the validity of features via domain knowledge, documentation via Large Language Models (LLMs) [233], or the detection of common issues – would benefit data creators to validate their datasets and assist users in selecting suitable data, enhancing NID research relevance, trustworthiness, and explainability.
7.4 Data Publishing
Formalizing data documentations: While reviewing the metadata of surveyed datasets, we observed that each was different. Although documentation of all scientifically-backed datasets follows a standard paper structure, some provide detailed data descriptions while others mention the data only briefly. This leads to under-documented datasets, reducing their credibility and usability. In addition, some datasets lack scientific backing or any documentation entirely.